OpenStudy (cutiecomittee123):

rewrite as an exponential equation ln(x+y)=5

OpenStudy (unknownunknown):

So, here you need to raise both sides by the exponential e. e^ln(x+y) = e^5. Can you see why we do this?

OpenStudy (cutiecomittee123):

Not really?? I know when you deal with logs that e is the base of ln

OpenStudy (cutiecomittee123):

I mean natural logs

OpenStudy (cutiecomittee123):

@unknownunknown

OpenStudy (unknownunknown):

That's right. So, you can think of ln(x) and e^x as being inverse to each other. \[e^{\ln x} = x \] Likewise, \[\ln e^{x} = x \] These are inverse functions. So, if we want to solve for x that's located in a natural logarithm, we need to apply its inverse function.

OpenStudy (cutiecomittee123):

Okay so now what

OpenStudy (cutiecomittee123):

@unknownunknown

OpenStudy (cutiecomittee123):

is that all?